3x – y = 28
3x + y = 14

Answer by jim_thompson5910(35256)

Hint:

Add the equations to get

3x – y = 28
3x + y = 14
------------
6x+0y = 42

6x = 42

solve 6x = 42 for x, then use this to find y.